OpenStudy (anonymous):

how long does it take a car going 65 mph to go 535 miles?

OpenStudy (anonymous):

t = s/v, where s = 535 miles = 860999.04 m v = 65 mph; so t = s/v = 860999.04 m/65 mph t = 552 days